# Empowering Embodied Manipulation: A Bimanual-Mobile Robot Manipulation Dataset for Household Tasks

### v1

@@ 4 Experiments
Rather than delving into an exhaustive analysis of different policy network design choices, the goal of our experiments is to explore the extent to which our dataset enhances the application of robot learning methods in executing household tasks. Our experiments are designed to address the following key questions: (a) Can BRMData effectively facilitate the learning of various robot manipulation methods? (b) Can BRMData effectively reflect a progressive increase in task difficulty? (c) Have we demonstrated fair and challenging testing methodologies to advance the research in embodied manipulation?
@@ 4.1 Experimental Setup
To investigate the performance and adaptability of various robot learning methods on BRMData, we conduct repetitive experiments using both single-task and multi-task models. The methods used in the single-task setting include Action Chunking with Transformers (ACT) [14] and Diffusion Policy (DP) [27]. The methods used in the multi-task experiments include MT-ACT [28], MT-ACT-EB3, and MT-ACT-R3M. Additional details can be found in Appendix E and F.
@@ 4.2 Experimental Results
4.2.1 Can BRMData facilitate the learning of various robot manipulation methods?
Table 3: Testing results of different robot learning models.
@@ 4.2.1 Can BRMData facilitate the learning of various robot manipulation methods?
To explore the performance of different robot learning methods, we select three complex robot manipulation tasks for testing, as shown in Table 3. In single-task experiments, the ACT method generally outperforms the DP method in terms of execution efficiency. Despite a lower SR in the Multiple Fruits Pick task, ACT achieves a superior MES of 3.61, indicating better efficiency. In the Cup Place task, ACT excels with a 90% SR and the highest MES of 6.06, highlighting its precision and speed in task execution. In multi-task experiments, MT-ACT proves to be highly adaptable and efficient, achieving a 90% SR and an MES of 4.78 in the Multiple Fruits Pick task, along with strong performance across other tasks. The experimental results demonstrate the effectiveness of BRMData in supporting the learning of various robot manipulation methods.
4.2.2 Can BRMData effectively reflect a progressive increase in task difficulty?
@@
Impact of Dual-Arm Coordination on Task Difficulty: As illustrated in Figure 4, the experimental results of single-task methods demonstrate varying degrees of success and efficiency, indicating the diverse difficulty levels in robot manipulation tasks. In the Bottle Pick task, both ACT and DP methods achieve high performance, with DP showing a SR of 80% and a MES of 7.84. However, when the task complexity increases to the dual-arm Bottle Handoff task, both SR and MES decline significantly for both methods. ACT’s SR drops to 75% and MES to 3.48, while DP’s SR falls to 60% and MES to 2.67. A similar trend is observed in fruit manipulation tasks. For Single Fruit Pick, ACT achieves an SR of 90% and a MES of 5.58, but in the dual-arm Multiple Fruits Pick task, ACT’s SR reduces to 60% and MES to 3.61, underscoring the difficulty in managing multiple items simultaneously. In placement tasks, Plate Place involves simpler, single-arm action, where DP maintains a high SR of 100% and a MES of 8.70. In contrast, the dual-arm Cup Place task results in lower performance, with DP’s SR at 80% and MES at 4.20. These comparisons clearly demonstrate that dual-arm tasks introduce significant complexity, leading to reduced success rates and efficiency.
Table 4: Testing results in robot manipulation tasks with human-robot interaction.
Method Single Fruit Pick Fruit Handover High Five
SR MES SR MES SR MES
ACT 90 5.58 85 8.92 100 10.28
DP 50 2.49 40 3.56 80 4.86
Impact of Human-Robot Interaction on Task Complexity: As illustrated in Table 4, human-robot interaction significantly increases task complexity. In the Fruit Handover task, DP’s SR drops to 40% compared to 50% in the Single Fruit Pick task. The coordination required for transferring objects between the robot and human demands precise timing and spatial alignment. Conversely, the High Five task, despite involving human-robot interaction, has a simpler dynamic without object manipulation, allowing for higher SR values. These observations highlight that the complexity of human-robot interaction depends heavily on the nature of the task, particularly whether it involves object manipulation.
Method Wine Wipe Garbage Recycle
@@
ACT 100 5.06 55 2.62
DP 80 3.38 40 1.57
Impact of Mobility on Task Complexity: The experimental results in Table 5 demonstrate that tasks requiring mobility are significantly more complex than stationary tasks. In the Wine Wipe task, which involves stationary manipulation, ACT achieves a perfect SR of 100% and an MES of 5.06, while DP achieves an SR of 80% and an MES of 3.38. However, when the task requires mobility, as in Garbage Recycle, performance drops notably: ACT’s SR decreases to 55% and MES to 2.62, and DP’s SR decreases to 40% and MES to 1.57. The dual requirement for seamless integration of navigation and manipulation in mobile tasks presents additional challenges, such as maintaining stability while moving and accurately reaching designated locations. These results underscore the challenges in maintaining high performance in tasks that involve mobility, thus demonstrating the robustness and utility of BRMData in evaluating the capabilities of robot learning methods under varied and complex conditions.
4.2.3 Have we demonstrated fair and challenging testing methodologies to advance the research in embodied manipulation?
@@
Fairness and Effectiveness of MES: Table 6 shows that the MES metric offers a comprehensive and fair measure of performance. By incorporating both SR and execution time, MES effectively penalizes methods that take longer to complete tasks, ensuring a balanced evaluation. For example, in the Multiple Fruits Pick task, despite ACT having a lower SR (60%) compared to DP (80%), ACT’s MES of 3.61 is higher than DP’s 2.94. This highlights ACT’s greater efficiency, demonstrating that MES can fairly assess the trade-offs between success rates and execution efficiency. Such a balance is crucial for real-world applications where both effectiveness and efficiency are essential. The MES also excels in highlighting the trade-offs between different approaches. For instance, in the High Five task, ACT achieves a high MES of 10.28, indicating both a high success rate and efficient execution. Conversely, DP achieves a lower MES of 4.86, pointing to longer execution times despite a reasonable SR of 80%. This shows that MES can effectively differentiate between methods based on both their accuracy and operational efficiency. Furthermore, MES helps to identify inefficiencies in more complex tasks, such as Garbage Recycle, where the need for both manipulation and mobility challenges the methods, resulting in lower MES values for both ACT (2.62) and DP (1.57). To further assess the robustness of imitation learning methods under varying conditions, additional experiments are conducted as detailed in Appendix G.
@@ 5 Conclusion
In summary, the BRMData dataset represents a significant advancement in the field of Embodied AI by addressing the critical shortage of comprehensive datasets for dual-arm mobile robot manipulation. By encompassing a range of bimanual and mobile manipulations, BRMData facilitates the exploration of complex interactions within dynamic environments that closely mimic real-world scenarios. This dataset not only supports the deployment of humanoid robots with enhanced capabilities but also provides a benchmark for testing and improving state-of-the-art models in robot manipulation. The diverse task configurations included in BRMData have proven instrumental in evaluating the adaptability and effectiveness of various robotic models. This has allowed for a detailed assessment of how different approaches perform under complex and varied conditions, thus contributing valuable insights into the development of more versatile and capable robotic systems. This work underscores the importance of such datasets in advancing our understanding of robotic capabilities and limitations, highlighting key areas for future exploration and development in the realm of embodied robots.
@@ D.2 Task Definitions
Appendix E Experimental Setup
@@
Appendix F Robot Learning Methods
F.1 Single-task Robot Learning Methods
@@
F.2 Multi-task Robot Learning Methods
@@
Table 8: Language Commands for Multi-task Methods
@@
Appendix G Robustness Testing for Robot Manipulation Tasks
@@
The repetitive robot manipulation experiments focus on evaluating the model’s inference execution and consistency, achieving high SRs due to the low-noise task environment and minimal background changes. To assess the robustness of robot learning methods, additional experiments are conducted by introducing significant changes to the tabletop environment and adding distractors. These changes aim to test how well the models maintain performance in the face of unexpected variations and increased complexity. The results, shown in Table 9, indicate a significant drop in both SR and MES compared to the initial repetitive tests. For instance, the ACT method’s SR for Plate Place drops from 100% to 40%, with a corresponding decrease in MES to 4.29. Similarly, DP’s performance in the Single Fruit Pick task falls to a 0% SR, indicating its inability to handle the increased complexity introduced by background changes and distractors. These findings underscore the critical importance of robustness testing in advancing the research in embodied manipulation. High performance in controlled, low-noise environments does not necessarily translate to real-world efficacy. The robustness tests reveal the limitations of the current models, particularly their struggles with environmental variability and unforeseen obstacles. Such testing methodologies are essential to understanding and improving the adaptability and resilience of robot manipulation models. Ultimately, the goal is to develop models that are not only precise and consistent in controlled settings but also resilient and versatile in diverse and challenging real-world environments. This highlights the ongoing need for refinement and enhancement of robot manipulation algorithms to achieve reliable performance in practical applications.
Appendix H Limitations and Future Work
@@
Appendix I Implementation Details
In this paper, the hyperparameters of the evaluated methods for the proposed dataset are shown in Table 10-11, respectively. Meanwhile, in the multi-task robot learning methods, we adopt the CLIP [35] network to encode the language commands. Besides, all the models are trained in a server equipped with 8 NVIDIA L40S graphics cards, an Intel Platinum 8468V CPU, and 1024GB of RAM.

### v2

@@ 4 Experiments
Rather than delving into an exhaustive analysis of different policy network design choices, the goal of our experiments is to explore the extent to which our dataset evaluates and enhances the application of robot manipulation learning methods in executing household tasks. Our experiments are designed to address the following key questions: (a) Is BRMData effective for testing various robot manipulation learning methods? (b) How do the properties of BRMData support the evaluation of methods?
@@ 4.1 Experimental Setup
To investigate the performance and adaptability of various robot manipulation learning methods on BRMData, we conduct repetitive experiments using both single-task and multi-task methods. The methods used in the single-task setting include Action Chunking with Transformers (ACT) [16] and Diffusion Policy (DP) [27]. The methods used in the multi-task experiments include MT-ACT [31], MT-ACT-EB3, and MT-ACT-R3M. Additional details of the experimental setup and the methods can be found in Appendix F and G.
@@ 4.2 Experimental Results
4.2.1 Is BRMData effective for testing various robot manipulation learning methods?
Table 3: Testing results of different robot manipulation learning methods.
@@ 4.2.1 Is BRMData effective for testing various robot manipulation learning methods?
To effectively test various robot manipulation learning methods on BRMData, we select three representative robot manipulation tasks: Multiple Fruits Pick (tabletop dual-arm manipulation), Cup Place (tabletop dual-arm interaction manipulation), Garbage Recycle (mobile manipulation). The testing results are shown in Table 3. In single-task method experiments, the ACT method generally outperforms the DP method in terms of execution efficiency. Despite a lower SR in the Multiple Fruits Pick task, ACT achieves a superior MES of 3.61, indicating better efficiency. In the Cup Place task, ACT excels with a 90% SR and the highest MES of 6.06, highlighting its precision and speed in task execution. This is because the denoising process of DP is time-consuming. Meanwhile, the experiments find that DP is good at diffusion learning of action sequences, but is not particularly sensitive to visual states, compared with ACT. Moreover, the generalization of both ACT and DP is weak, particularly when there is a perturbation in the object’s position, which often results in an inability to complete the task.
In multi-task method experiments, MT-ACT proves to be highly adaptable and efficient, achieving a 90% SR and an MES of 4.78 in the Multiple Fruits Pick task, along with strong performance across other tasks. On the contrary, MT-ACT-R3M uses a pre-trained R3M model but still performs poorly because it is limited by the smaller image size input, resulting in the loss of visual information. Compared to MT-ACT, MT-ACT-EB3 loses visual information, has larger model parameters, and requires longer inference time. The experimental results show that different robot manipulation methods can be tested on our dataset. The possible users can test both single-task learning methods and multi-task learning methods and make a comparison with the existing methods. Besides, to further assess the robustness of the learning methods under varying conditions, additional experiments are conducted as detailed in Appendix I.
4.2.2 How do the properties of BRMData support the evaluation of methods?
The tasks of BRMData have the properties of task diversity, progressive difficulty, and sensory data comprehensiveness. They can facilitate the verification of method effectiveness across diverse tasks. Both the basic capability to finish simple tasks and the upper boundary for complex tasks of the methods can be illustrated.
@@ 4.2.2 How do the properties of BRMData support the evaluation of methods?
Verification of method effectiveness on diverse tasks. As illustrated in Figure 4, the experimental outcomes from single-task methodologies, namely ACT and DP, exhibit a spectrum of success and operational efficiency, thereby offering a robust validation of the evaluative methods employed. Within the ambit of single-arm operations, such as Bottle Pick, Single Fruit Pick, and Bottle Place, the ACT approach registers an SR between 90% and 100%, and an MES ranging from 5.58 to 11.04. Conversely, the DP strategy secures an SR of 50% to 90%, with MES values spanning 2.49 to 8.7. In dual-arm tasks, the ACT method attains a SR of 60% to 90%, alongside an MES between 3.48 and 6.06, whereas DP records a SR of 60% to 80%, and an MES ranging from 2.94 to 4.2. The testing results demonstrate that the ACT method leads to a relatively higher success rate and higher efficiency. The application across a varied set of tasks on BRMData substantiates the effectiveness of the manipulation methods under investigation.
Table 4: Testing results in robot manipulation tasks without and with human-robot interaction.
Method Single Fruit Pick Fruit Handover
@@
ACT 90 5.58 85 8.92
DP 50 2.49 40 3.56
Illustration of method capability boundaries on tasks with progressive difficulty. As illustrated in Table 4, the tasks of Single Fruit Pick and Fruit Handover involve the picking and placing of a single fruit, differing only in the occurrence of human-robot interaction. ACT method achieves an SR of 90% and an MES of 5.58, whereas the DP method records an SR of 50% and a MES of 2.49. These results indicate that the ACT method performs effectively in basic tasks, while the DP method exhibits moderate performance. In the more challenging Fruit Handover task, the performance of both methods declines, with ACT achieving an 85% SR and an 8.92 MES, and DP a 40% SR and a 3.56 MES. To be noticed, the task Single Fruit Pick involves repeating the picking and placing twice time, thus resulting in a small MES. This comparative analysis suggests that the ACT method not only maintains superior task execution in complex scenarios but also hints at a potentially higher threshold of capability. Conversely, the DP method’s diminished performance in the Fruit Handover task suggests it may have reached its capability limit for tasks of increased difficulty. In addition, the impact of mobility on task complexity is discussed in Appendix H.
Effectiveness analysis of the MES metric for evaluating methods. Table 5 shows that the MES metric offers a comprehensive and fair measure of performance. By incorporating both SR and execution time, MES effectively penalizes methods that take longer to complete tasks, ensuring a balanced evaluation. For example, in the Multiple Fruits Pick task, despite ACT having a lower SR (60%) compared to DP (80%), ACT’s MES of 3.61 is higher than DP’s 2.94. This highlights ACT’s greater efficiency, demonstrating that MES can fairly assess the trade-offs between success rates and execution efficiency. Such a balance is crucial for real-world applications where both effectiveness and efficiency are essential. The MES also excels in highlighting the trade-offs between different approaches. For instance, in the High Five task, ACT achieves a high MES of 10.28, indicating both a high success rate and efficient execution. Conversely, DP achieves a lower MES of 4.86, pointing to longer execution times despite a reasonable SR of 80%. This shows that MES can effectively differentiate between methods based on both their accuracy and operational efficiency. Furthermore, MES helps to identify inefficiencies in more complex tasks, such as Garbage Recycle, where the need for both manipulation and mobility challenges the methods, resulting in lower MES values for both ACT (2.62) and DP (1.57).
@@ 5 Conclusion
In this paper, BRMData is proposed to address the critical shortage of comprehensive datasets for dual-arm mobile robot manipulation, which represents a significant advancement in the field of embodied manipulation. By encompassing a range of bimanual and mobile manipulations, BRMData facilitates the exploration of complex interactions within dynamic environments that closely mimic real-world scenarios. This dataset not only supports the deployment of humanoid robots with enhanced capabilities but also provides a benchmark for testing and improving state-of-the-art models in robot manipulation. The diverse task configurations included in BRMData have proven instrumental in evaluating the adaptability and effectiveness of various robot manipulation learning methods. This has allowed for a detailed assessment of how different approaches perform under complex and varied conditions, thus contributing valuable insights into the development of more versatile and capable robotic systems. This work underscores the importance of such datasets in advancing our understanding of robotic capabilities and limitations, highlighting key areas for future exploration and development in the realm of embodied robots.
@@ D.2 Task Definitions
Appendix E Robot Platform
Figure 15: Illustration of robot platform.
Appendix F Experimental Setup
@@
Appendix G Robot Manipulation Learning Methods
G.1 Single-task Robot Manipulation Learning Methods
@@
G.2 Multi-task Robot Manipulation Learning Methods
@@
Table 7: Language commands for multi-task learning methods
@@
Appendix H Impact of Mobility on Task Complexity
Method Wine Wipe Garbage Recycle
SR MES SR MES
ACT 100 5.06 55 2.62
DP 80 3.38 40 1.57
The experimental results in Table 8 demonstrate that tasks requiring mobility are significantly more complex than stationary tasks. In the Wine Wipe task, which involves stationary manipulation, ACT achieves a perfect SR of 100% and an MES of 5.06, while DP achieves an SR of 80% and an MES of 3.38. However, when the task requires mobility, as in Garbage Recycle, performance drops notably: ACT’s SR decreases to 55% and MES to 2.62, and DP’s SR decreases to 40% and MES to 1.57. The dual requirement for seamless integration of navigation and manipulation in mobile tasks presents additional challenges, such as maintaining stability while moving and accurately reaching designated locations. These results underscore the challenges in maintaining high performance in tasks that involve mobility, thus demonstrating the robustness and utility of BRMData in evaluating the capabilities of robot learning methods under varied and complex conditions.
Appendix I Robustness Testing for Robot Manipulation Tasks
@@
The repetitive robot manipulation experiments focus on evaluating the model’s inference execution and consistency, achieving high SRs due to the low-noise task environment and minimal background changes. To assess the robustness of robot manipulation learning methods, additional experiments are conducted by introducing significant changes to the tabletop environment and adding distractors. These changes aim to test how well the models maintain performance in the face of unexpected variations and increased complexity. The results, shown in Table 9, indicate a significant drop in both SR and MES compared to the initial repetitive tests. For instance, the ACT method’s SR for Plate Place drops from 100% to 40%, with a corresponding decrease in MES to 4.29. Similarly, DP’s performance in the Single Fruit Pick task falls to a 0% SR, indicating its inability to handle the increased complexity introduced by background changes and distractors. These findings underscore the critical importance of robustness testing in advancing the research in embodied manipulation. High performance in controlled, low-noise environments does not necessarily translate to real-world efficacy. The robustness tests reveal the limitations of the current models, particularly their struggles with environmental variability and unforeseen obstacles. Such testing methodologies are essential to understanding and improving the adaptability and resilience of robot manipulation models. Ultimately, the goal is to develop models that are not only precise and consistent in controlled settings but also resilient and versatile in diverse and challenging real-world environments. This highlights the ongoing need for refinement and enhancement of robot manipulation algorithms to achieve reliable performance in practical applications.
Appendix J Limitations and Future Work
@@
Appendix K Implementation Details
In this paper, the hyperparameters of the evaluated methods for the proposed dataset are shown in Table 10-11, respectively. Meanwhile, in the multi-task robot manipulation learning methods, we adopt the CLIP [37] network to encode the language commands. Besides, all the models are trained in a server equipped with 8 NVIDIA L40S graphics cards, an Intel Platinum 8468V CPU, and 1024GB of RAM.